OpenStudy (anonymous):

1.calculate the amount of heat produced in an electric heater of resistance 1000 ohms if 6A of current is passed through it for 10 minutes. 2.400J of heat is produced in 4s in a 4 ohm resistor.Find potential difference across the resistor. 3. Calculate the power of a source if it produces an energy of 750J in 30s. 4. Which uses more energy , a 500W TV set in 1 hour or a 1200W toaster in 10 minutes. 5. Calculate the electrical energy consumed by a 1200W toaster in 10 minutes.

OpenStudy (anonymous):

5. Energy (J) = Power (W) * time (s)

OpenStudy (anonymous):

4. use this ^

OpenStudy (anonymous):

3. 750 = P * (30/3600) solve for P

OpenStudy (anonymous):

1. Not sure but maybe: V=IR V = 1000*6 = 6000v P = IV = 6000*1000 Energy = Power * Time Energy = 6000kW * (10/60)h =1000kWh

OpenStudy (anonymous):

1.\[E=I ^{2}Rt\] where you put \[I=6A\]\[R= 6Ohm\]\[t=10*60\] (where t is in second unit) then you will get the amount of heat(E) produced.

OpenStudy (anonymous):

thank you so much psujono and biswajit_paul.

OpenStudy (anonymous):

how do I solve the second question?

OpenStudy (anonymous):

2.400J of heat is produced in 4s in a 4 ohm resistor.Find potential difference across the resistor. E=I^2*R*t 400 = I^2*4*4 25=I^2 5=I V=IR = 5*4=20 volts

OpenStudy (anonymous):

thank you again psujono

OpenStudy (anonymous):

No probs

OpenStudy (anonymous):

so which one is the correct answer for the first question.Please guide me.

OpenStudy (shane_b):

1. Calculate the amount of heat produced in an electric heater of resistance 1000 ohms if 6A of current is passed through it for 10 minutes. \[E=Pt\]\[P=I^2R\]\[E=(6A)^2(1000 \Omega)(600s)=21.6MJ\] 2. 400J of heat is produced in 4s in a 4 ohm resistor.Find potential difference across the resistor. \[E=400J=I^2(4 \Omega)(4s)\]\[I=5A\]\[V=IR=(5A)(4 \Omega)=20V\] 3. Calculate the power of a source if it produces an energy of 750J in 30s. \[P=W=J/S=\frac{750J}{30s}=25W\] 4. Which uses more energy , a 500W TV set in 1 hour or a 1200W toaster in 10 minutes. \[E=Pt\]\[E_{tv}=(500W)(3600s)=1.8MJ\]\[E_{toaster}=(1200W)(600s)=720KJ\] 5. Calculate the electrical energy consumed by a 1200W toaster in 10 minutes. \[1200W*600s=720KJ\]

OpenStudy (anonymous):

thank you so much Shane_B.